箱线图显示了商店 A 和商店 B 售出的鸡蛋质量(克)的分布情况。

  1. (i)

    商店 A 售出的鸡蛋中,有 25 个鸡蛋的质量超过 59 克。

    求商店 A 售出的鸡蛋总数。

    答 ... 个鸡蛋

  2. (ii)

    对 A 商店和 B 商店出售的鸡蛋质量的平均值进行比较并发表评述,再对其质量的分布进行比较并发表评述。

    请用数据来支持你的回答。

    1. ...

    ...

    1. ...

    ...

  3. (8)
    1. (b)

      在 80 个鸡蛋的样本中,有 4 个是有裂缝的。

      1. (i)

        从样本中随机抽取一个鸡蛋。

        求该鸡蛋有裂纹的概率。

        答案 ...

      2. (ii)

        从样本中随机抽取两个鸡蛋。

        求两个鸡蛋都有裂纹的概率。

        答案 ...

      3. (iii)

        从样本中随机抽取三个鸡蛋。

        求至少有一个鸡蛋破裂的概率。

        答案 ...

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

8 (a)(i)

1. Relate masses above the upper quartile to one quarter of the total number of eggs.

For Shop A, Q3=59Q_3 = 59 g.

So the eggs with masses more than 5959 g make up 25%25\% of the total.

25% of total=2525\% \text{ of total} = 25

Total=25÷0.25=100\text{Total} = 25 \div 0.25 = 100

答案

100100 eggs

8 (a)(ii)

1. Compare the averages using the medians from the box-and-whisker plots.

For a box-and-whisker plot, the average is usually compared using the median.

Shop A median =54=54 g.

Shop B median =50=50 g.

Since 54>5054>50, Shop A has the higher average mass.

2. Compare the distributions using a measure of spread from the box-and-whisker plots.

Use the interquartile range to compare spread.

Shop A:
IQR=Q3−Q1=59−48=11\text{IQR}=Q_3-Q_1=59-48=11 g.

Shop B:
IQR=Q3−Q1=58−46=12\text{IQR}=Q_3-Q_1=58-46=12 g.

Since 12>1112>11, Shop B has a slightly greater spread in the middle 50%50\% of the masses.

答案
  1. The average mass of eggs in Shop A is higher than in Shop B, since the median for Shop A is 5454 g while the median for Shop B is 5050 g.

  2. The masses of eggs in Shop B are more spread out in the middle 50%50\% of the data, since its interquartile range is 58−46=1258-46=12 g, while Shop A's interquartile range is 59−48=1159-48=11 g.

8 (a)(8)(b)(i)

1. Find the probability that the selected egg is cracked.

There are 44 cracked eggs out of 8080 eggs.

P(cracked)=480=120P(\text{cracked})=\frac{4}{80}=\frac{1}{20}

答案

120\dfrac{1}{20}

8 (a)(8)(b)(ii)

1. Find the probability that the first egg selected is cracked.

There are 44 cracked eggs out of 8080 eggs.

P(first cracked)=480P(\text{first cracked})=\frac{4}{80}

2. Account for no replacement and multiply the probabilities.

After selecting one cracked egg, there are 33 cracked eggs left out of 7979 eggs.

P(both cracked)=480×379=126320=31580P(\text{both cracked})=\frac{4}{80}\times\frac{3}{79}=\frac{12}{6320}=\frac{3}{1580}

答案

1158\dfrac{1}{158}

8 (a)(8)(b)(iii)

1. Find the probability that none of the three selected eggs is cracked.

There are 80−4=7680-4=76 uncracked eggs.

P(none cracked)=7680×7579×7478P(\text{none cracked})=\frac{76}{80}\times\frac{75}{79}\times\frac{74}{78}

2. Use the complement to find the probability that at least one egg is cracked.

P(at least one cracked)=1−7680×7579×7478P(\text{at least one cracked})=1-\frac{76}{80}\times\frac{75}{79}\times\frac{74}{78}

=1−56157900=1-\frac{5615}{7900}

=22857900=229790=\frac{2285}{7900}=\frac{229}{790}

答案

229790\dfrac{229}{790}

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1对 A、B 两家工厂生产的两批各 400 根钢条的质量进行了测量和记录。下方的累积频数图像显示了这些质量。 对于工厂 A 生产的钢条,求 中位数。 答案 ... kg 四分位距。 答案 ... kg 工厂 B 生产的钢条比工厂 A 生产的钢…
  2. 2茎叶图显示了某班男生和女生的测试成绩。 [茎叶图印刷为两列,左列标为“男生”(Boys),右列标为“女生”(Girls),共用中间的一列茎。键(男生)8 | 3 表示 38。键(女生)4 | 1 表示 41。] 艾尔文在下方的箱线图中表示了…
  3. 3该累积频数图展示了某一天在餐厅 A 消费的 60 名顾客在接受服务前的等待时间(以分钟为单位)的分布。 利用该曲线估计: 中位数等待时间, 答:... min 四分位距, 答:... min 等待时间的第 35 百分位数。 答:... mi…
  4. 4两家电话服务店(A 店和 B 店)各 40 名顾客的等待时间 t(单位:分钟)如下。 计算 A 店的平均等待时间。 答:___ 分钟 解释为什么你对 (a)(i) 的回答只是一个估计值。 答 ... 计算 A 店的标准差。 答:___ 分钟…
  5. 5某一天班上 18 名女生的体温显示在下面的茎叶图中。 (茎叶图显示在下方并在 figure_description 中重现。) 键:37 | 1 表示 37.1∘C37.1^\circ\text{C} 已知极差为 1.8∘C1.8^\circ\text{C}…